# Reduction Complexities in Set Theory

### v4

@@ 3.2 Reductions to DecCard
Is there (provably in ZFC) a reduction of OrdCard to DecCard which provably in ZFC improves on the naive approach explained above in the strong sense that, for some $\alpha\in\text{On}$, the cardinality of the number of calls required on input $\xi>\alpha$ will be strictly smaller than that of $\xi$? By a construction similar to the above, we can see that, at least relative to the existence of a proper class of inaccessible cardinals in $L$, this is not the case. This is the point of the next theorem, for the preparation of which we prove to lemmas.
Lemma 4
Proof
@@
Assume that there is a proper class of inaccessible cardinals in LL. [^5] This follows from, but is much weaker than, the existence of $0^{\sharp}$. The it is consistent with ZFC that there is no ordinal $\alpha$ and no parameter-program $(P,\rho)$ that reduces OrdCard to DecCard and such that, for all $\beta>\alpha$, $(P,\rho)$ makes less than $\text{card}(\beta)$ many calls to DecCard in the input $\beta$.
@@
As explained above, we identify parameter-programs with ordinals. By assumption, for $\iota\in\text{On}$, there is a $\iota$-th inaccessible cardinal in $L$, which we denote by $\kappa_{\iota}$. We will avoid complications for our construction that would arise from inaccessible cardinals that are themselves limits of inaccessible cardinals and thus only consider successors in this sequence; accordingly, we write $\kappa(\alpha,\beta)$ for $\kappa_{p(\alpha,\beta)+1}$. Thus, we have reserved, for each parameter program $\alpha$, a proper class $C_{\alpha}$ of $L$-inaccessible cardinals (which are not limits or such cardinals), namely $C_{\alpha}:=\{\kappa(\alpha,\iota):\iota\in\text{On}\}$. We assume without loss of generality (e.g., by replacing $\kappa(\alpha,\beta)$ with $\kappa(\alpha,\beta+\alpha)$ that $\kappa(\alpha,\beta)>\text{max}(\alpha,\beta)$.
@@
We now again proceed by an iterated forcing, sabotaging, for each pair $(\gamma,\alpha)$ of a (coded) parameter-program $\gamma$ and an ordinal $\alpha$ the claim that, for $\delta>\alpha$, $\gamma$ achieves a reduction of OrdCard to DecCard in complexity $\text{card}(\delta)$. Noting that DecCard is easily definable as a class function, let us denote by $H^{M}:\text{On}\rightarrow\{0,1\}$ an (in fact, the unique) DecCard function in the transitive ZFC-model $M$; when the model is clear from the context, the superscript $M$ will be dropped. By replacing the parameter-program $\gamma$ by the parameter-program $\gamma^{\prime}$ guaranteed to exist by Lemma 4, we may assume without loss of generality that each parameter-program $\gamma$ we consider calls $H$ at most once on each ordinal.
We now explain one step of the iteration. Let an ordered pair $(\gamma,\alpha)$ of a parameter-program $\gamma$ and an ordinal $\alpha$ be given. Consider the cardinal $\kappa:=\kappa(\gamma,\alpha)$. Let us denote by $\tilde{H}$ the function that works like $H$, but returns 0 on input $\kappa$. [^6] Clearly, $\overline{H}$ is not a DecCard function for any model of set theory, as it claims that a limit of cardinals can be a non-cardinal. The idea is that, in the construction below, we will collapse $\kappa$ and leave all relevant instances of $\overline{H}$ untouched (where “relevant” means: sufficient to ensure that $\gamma^{H}(\kappa)$ does not yield the right result in less than $\text{card}(\kappa)$ many steps). We assume that our construction has already taken care of all earlier pairs in the natural well-ordering of pairs of ordinals and that we have thus obtained a transitive $M\models\text{ZFC}$, which has the following property: If $\nu$ is the supremum of the $L$-inaccessible cardinals strictly smaller than $\kappa$, then, from $(\mu+)^{M}$ upwards, the $M$-cardinals are the same as the $L$-cardinals. Our construction will ensure that (i) cardinals and cofinalities up to (and including) $(\mu+)^{M}$ are not changed, (ii) cardinals greater than or equal to $(\kappa^{+})^{M}$ are not changed (and thus remain the same as in $L$), and in particular regular cardinals greater than or equal to $(\kappa^{+})^{M}$ remain regular, (iii) $\kappa$ is collapsed (so that, in the generic extension $M[G]$, $\tilde{H}(\kappa)$ will agree with $H^{M[G]}(\kappa)$) and (iv) in the extension $M[G]$, GCH holds from $(\kappa^{+})^{M[G]}=(\kappa^{+})^{M}$, i.e., for every cardinal $\nu\geq\kappa^{+}$, we have $2^{\nu}=\nu^{+}$ in $M[G]$.
We define the following class functions from (pairs of) ordinals to ordinals. For all $\gamma,\alpha,\tau\in\text{On}$, let $s(\gamma,\alpha,\tau)$ be the supremum of ordinals $\alpha$ for which the computation of $\gamma^{\tilde{H}}(\alpha)$ makes requests to $\tilde{H}$ up to time $\tau$. We let $F_{\gamma}(\alpha,\beta)$ be the minimal ordinal $\tau$ such that, at time $\tau$ in the computation of $\gamma^{\tilde{H}}(\alpha)$, we have $s(\gamma,\alpha)\geq\beta$. Finally, we let $G_{\gamma}(\alpha,\beta)$ be the order type of times in the computation of $\gamma^{\tilde{H}}(\alpha)$ up to time $F_{\gamma}(\alpha,\beta)$ at which calls to $\tilde{H}$ concerning ordinals $\alpha$ were made.
Claim
@@
    The function $\beta\mapsto F_{\gamma}(\alpha,\beta)$ is continuous.
@@
    The function $\beta\mapsto G_{\gamma}(\alpha,\beta)$ is continuous.
@@
    If $\kappa>\omega$ is regular and the image of $\kappa$ under the function $\beta\mapsto G_{\gamma}(\kappa,\beta)$ is unbounded in $\kappa$, then this function has a club of fixed-points in $\kappa$. [^7] This is of course very similar to the fixed-points theorem for normal functions (see, e.g., [13], Exercise 8.1) – and the proof is a simple adaptation of the proof for that theorem – but not a direct application, as the function $\beta\mapsto G_{\gamma}(\kappa,\beta)$ need not be strictly increasing.
Proof
By definition, it is clear that $\beta\mapsto F_{\gamma}(\alpha,\beta)$ and $\beta\mapsto G_{\gamma}(\alpha,\beta)$ are both non-decreasing.
  1. 1.
    Let $(\beta_{\iota}:\iota<\lambda)$ be an increasing sequence of ordinals in $\kappa$ of limit length $\lambda<\kappa$, and let $\beta:=\sup_{\iota<\lambda}\beta_{\iota}$.
    • $\leq$”: Let $\tau^{\prime}<F_{\gamma}(\alpha,\beta)$. Thus, at time $\tau^{\prime}$, the ordinal $\beta$ was not attained as the supremum of the accessed ordinals. By definition of $\beta$, there is $\xi<\lambda$ such that $\beta_{\xi}>s(\gamma,\alpha,\tau^{\prime})$. So $F_{\gamma}(\alpha,\beta_{\xi})>\tau^{\prime}$. Because $\sup_{\iota<\lambda}F_{\gamma}(\alpha,\beta_{\iota})\geq F_{\gamma}(\alpha,\beta_{\xi})>\tau^{\prime}$, it follows that $\sup_{\iota<\lambda}F_{\gamma}(\alpha,\beta_{\iota})>\tau^{\prime}$.
    • $\geq$”: Let $\tau^{\prime}<\sup_{\iota<\lambda}F_{\gamma}(\alpha,\beta_{\iota})$, and let $\xi<\lambda$ be such that $\tau^{\prime}<F_{\gamma}(\alpha,\beta_{\xi})$. Thus $s(\gamma,\alpha,\tau^{\prime})<\beta_{\xi}<\beta$, i.e., at time $\tau^{\prime}$, $\beta$ was not yet attained as the supremum of the accessed ordinals. So $F_{\gamma}(\alpha,\beta)>\tau^{\prime}$.
  2. 2.
    Again, let $(\beta_{\iota}:\iota<\lambda)$ be an increasing sequence in $\alpha$ with $\beta:=\sup_{\iota<\lambda}\beta_{\iota}<\alpha$. By (1), we have $F_{\gamma}(\alpha,\beta)=\sup_{\iota<\lambda}F_{\gamma}(\alpha,\beta_{\iota})$. Now, for all $\iota<\xi<\lambda$, the sequence of times before $\xi$ at which calls to $\tilde{H}$ concerning ordinals $\alpha$ were made is an end-extension of the corresponding sequence up to time $\iota$. Thus, that sequence up to time $\beta$ is just the (increasing) union of these sequences up to the times $\beta_{\xi}$ with $\xi<\lambda$, and hence the corresponding order type at time $\beta$ is the supremum of the sequence of the order types at times $\beta_{\xi}$ for $\xi<\lambda$.
  3. 3.
    Let $C$ denote the set of fixed points, i.e. $C:=\{\iota<\kappa:G_{\gamma}(\kappa,\iota)=\iota\}$.
    • $C$ is closed: Let $(\xi_{\iota}:\iota<\lambda)\subseteq C$ be a sequence with $\xi:=\sup_{\iota<\lambda}\xi_{\iota}<\kappa$. Then $G_{\gamma}(\alpha,\xi)=G_{\gamma}(\alpha,\sup_{\iota<\lambda}\xi_{\iota})=\sup_{\iota<\lambda}G_{\gamma}(\alpha,\xi_{\iota})=\sup_{\iota<\lambda}\xi_{\iota}=\xi$, where the second equality uses (2) and the third uses the definition of $C$.
    • $C$ is unbounded: Let $\lambda<\kappa$. By the unboundedness assumption, there is $\iota_{0}<\kappa$ such that $G_{\gamma}(\alpha,\iota_{0})>\lambda$. Define recursively $\iota_{k+1}:=\text{min}\{\iota<\kappa:G_{\gamma}(\alpha,\iota)>\text{max}\{G_{\gamma}(\alpha,\iota_{k}),\iota_{k}\}\wedge\iota>G_{\gamma}(\alpha,\iota_{k})\}$. Again, the unboundedness assumption yields that $\iota_{k}$ will exist for all $k\in\omega$. It follows by definition that, for all $k\omega$, we have $G_{\gamma}(\alpha,\iota_{k})<\iota_{k+1}$ and $\iota_{k}<G_{\gamma}(\alpha,\iota_{k+1})$. More concisely, we have $G_{\gamma}(\alpha,\iota_{k})<\iota_{k+1}<G_{\gamma}(\alpha,\iota_{k+2})$ for all $k\in\omega$, so $\sup_{k\in\omega}G_{\gamma}(\alpha,\iota_{k})=\sup_{k\in\omega}\iota_{k}$. Let $\varphi:=\sup_{k\in\omega}\iota_{k}$. Then $\varphi>\lambda$, and moreover, we have $G_{\gamma}(\alpha,\varphi)=G_{\gamma}(\alpha,\sup_{k\in\omega}\iota_{k})=\sup_{k\in\omega}G_{\gamma}(\alpha,\iota_{k})=\sup_{k\in\omega}\iota_{k}=\varphi$, so $\varphi$ is also a fixed point.
Let us consider the first $\kappa$ calls that $\gamma^{\overline{H}}(\kappa)$ makes to $\overline{H}$. Denote by $S$ the set of ordinals to which these calls are made. By Lemma 5, these requests all concern ordinals strictly less than $\kappa^{+}$ (for that lemma ensures that at least $\kappa^{+}$ many calls need to be made before the first call to an ordinal $\geq\kappa^{+}$ can be made). [^8] Strictly speaking, Lemma 5 is not directly applicable, as $\tilde{H}$ is not a DecCard function; however, the proof goes through for $\tilde{H}$ as well.
We now distinguish two cases:
@@
$S\cap\kappa$ is bounded in $\kappa$ by some ordinal $\delta<\kappa$. By making $\delta$ larger if necessary, we can assume without loss of generality that $\delta>(\mu^{+})^{M}$.
Proof
In this case, if the computation does not halt, or halts and outputs an ordinal which is not an $L$-cardinal strictly below $\kappa$, let $\lambda\in(\delta,\kappa)$ be an arbitrary $L$-cardinal. Otherwise, the output will be some $L$-cardinal strictly less than $\kappa$, which we denote by $\lambda$.
Let $\eta:=\text{max}(\lambda,\delta)$. Then $\eta<\kappa$. By inaccessibility of $\kappa$ in $L$ (and thus, by (ii), in $M$, since $\kappa>(\mu^{+})^{M}$), we have $(\eta^{+})^{M}=(\eta^{+})^{L}<\kappa$. Force with $\text{Coll}(\eta^{+},\kappa)$. Let $M[G]$ be the generic extension. This forcing is $\eta^{+}$-closed; hence, for all $\nu\in S$, we have that $\nu$ is a cardinal in $M$ if and only if it is a cardinal in $M[G]$. Consequently, we have $\tilde{H}(\nu)=H^{M[G]}(\nu)$ for all $\nu\in S$. Moreover, the forcing collapses $\kappa$, so $\kappa$ is not a cardinal in $M[G]$. Thus, we also have $\tilde{H}(\kappa)=H^{M[G]}(\kappa)$. Finally, the ordinals in $(\kappa,(\kappa^{+})^{M})$ are not cardinals both in $M$ and in $M[G]$. It follows that $\tilde{H}(\nu)=H^{M[G]}(\nu)$ for all ordinals for which $\tilde{H}$ is called in the computation of $\gamma^{\tilde{H}}(\kappa)$.
Note that, by the case assumption, the computations $\gamma^{\tilde{H}}(\kappa)$ and $\gamma^{H^{M[G]}}(\kappa)$ are actually the same. Thus, if $\gamma^{\tilde{H}}(\kappa)$ diverges, then so does $\gamma^{H^{M[G]}}(\kappa)$, and if it converges, then both computations have the same output.
If $\gamma^{\overline{H}}(\kappa)\neq\kappa$, then that output is $\lambda$, and if $\gamma^{\overline{H}}(\kappa)=\kappa$, the output will be $\kappa$. However, in $M[G]$, we have $\kappa>\text{card}(\kappa)=\eta^{+}>\lambda$, so in both cases, the output will not be the cardinality of $\kappa$ in $M[G]$. Thus $\gamma$ does not reduce OrdCard to DecCard in $M[G]$.
We need to check that (i)-(iv) from above are satisfied. We already did (iii). By $\eta^{+}$-closure, all cardinals and cofinalities less than or equal to $\eta$ are preserved, which includes all cardinals and cofinalities less than or equal to $(\mu^{+})^{M}$; hence (i) is satisfied. Moreover, the forcing $\text{Coll}(\eta^{+},\kappa)$ has cardinality $\kappa$ [^9] We briefly recall the (standard) argument for the sake of the reader: The forcing has size $\kappa^{\eta}$, and since $\kappa$ is regular in $M$ by (ii) and $\eta<\kappa$, any function from $\eta$ to $\kappa$ is bounded in $\kappa$. Thus $\kappa^{\eta}=\Sigma_{\nu<\kappa}\nu^{\eta}$; since, by (iv), GCH holds below $\kappa$ from some $\theta<\kappa$ on, we have $\Sigma_{\nu<\kappa}\nu^{\eta}=\Sigma_{\theta<\nu<\kappa}\nu^{\eta}=\Sigma_{\theta<\nu<\kappa}\nu^{+}=\Sigma_{\theta^{+}<\nu<\kappa}\nu=\kappa$. and thus trivially satisfies the $\kappa$-cc. [^10] This could also be deduced from [15], Lemma 4.4, p. 264. It follows that (regular) cardinals from $\kappa^{+}$ upwards are preserved (see, e.g., [15], Corollary 6.11), so we have (ii) in $M[G]$. Finally, by the $\kappa$-$c$, $\text{Coll}(\eta^{+},\kappa)$ has at most $\kappa^{\kappa}\leq(2^{\kappa})^{\kappa}=2^{\kappa\cdot\kappa}=2^{\kappa}$ many antichains in $M$, which, by (iv) in $M$, is equal to $\kappa^{+}$. Thus, for an $M$-cardinal $\theta\geq\kappa^{+}$, there are at most $(2^{\kappa})^{\theta}=2^{\kappa\cdot\theta}=2^{\theta}=\theta^{+}$ many nice names for subsets of $\theta$ (where the last step again uses (iv) in $M$), and so $(2^{\theta})^{M[G]}=(2^{\theta})^{M}=\theta^{+}$ for $\theta\geq\kappa^{+}$, establishing (iv). [^11] Cf., e.g., [15], proof of Theorem 6.17, for this standard argument.
Note that, by (i), further extensions will not alter any of the cardinals relevant for the computation $\gamma^{\tilde{H}}(\kappa)$, so the result of this computation remain the same (and wrong) in all subsequent extensions.
@@
$S\cap\kappa$ is unbounded in $\kappa$.
Proof
By our assumption on $\gamma$ that no ordinal is requested twice, the number of requests to $\tilde{H}$ that concern ordinals below a given ordinal $\xi$ is always strictly less than $(\xi^{+})^{M}$. Let $g$ be the restriction of the (class) function $\beta\mapsto G_{\gamma}(\kappa,\beta)$ to $\kappa$. It follows by inaccessibility of $\kappa$ that $g$ is actually a function from $\kappa$ to $\kappa$. By part (3) of our claim above, there is a club $C\subseteq\kappa$ of fixed points of $g$. Moreover, by inaccessibility of $\kappa$, the set of $M$-cardinals below $\kappa$ is also a club. Thus, the intersection of these two clubs is a club. Pick an element $\lambda$ of this club which is larger than $(\mu^{+})^{M}$. Let $\tau:=F_{\gamma}(\kappa,\lambda)$; thus, before time $\tau$, all calls to $\tilde{H}$ in the computation of $\gamma^{\tilde{H}}(\kappa)$ that concerned ordinals $\kappa$ actually concerned ordinals $\lambda$.
Now force with $\text{Coll}(\lambda,\kappa)$. Let $M[G]$ be the generic extension. In $M[G]$, we have $\text{card}(\kappa)=\lambda$. The forcing is $\lambda$-closed and thus preserves $\lambda$ and all lower cardinals; since $\lambda>(\mu^{+})^{M}$, this implies (i). (ii) and (iv) follows from the $\kappa$-cc as in the first case, and (iii) also follows as in the first case. Moreover, in $M[G]$, we have $\text{card}(\kappa)=\lambda$.
We claim that the computations $\gamma^{\tilde{H}}(\kappa)$ and $\gamma^{H^{M[G]}}(\kappa)$ agree up to time $\tau$. Since at least [^12] Note that there could be additional requests to ordinals in $\kappa,\kappa^{+})$. $\lambda$ many requests to $\tilde{H}$ were made up to this point (by definition of $g$), and $\lambda=\text{card}(\kappa)$, we have that $\gamma^{H^{M[G]}}(\kappa)$ makes at least $\text{card}^{M[G]}(\kappa)$ many calls to $H$ in $M[G]$, which is what we wanted. Also note that, by (i), all cardinals relevant to this fact will be preserved in all further extensions.
To see the claim, note first that, since cardinals less than or equal to $\lambda$ are preserved, we have $\tilde{H}(\xi)=H^{M[G]}(\xi)$ for all $\xi<\lambda$. By (iii), we have $\tilde{H}(\kappa)=H^{M[G]}(\kappa)=0$. Moreover, as in case $1$, we have $\tilde{H}(\xi)=H^{M[G]}(\xi)=0$ for all $\xi\in(\kappa,\kappa^{+})$. By assumption, $\gamma^{\tilde{H}}(\kappa)$ makes no calls to ordinals $\xi\in[\lambda,\kappa)$ up to time $\tau$; if $\gamma^{H^{M[G]}}(\kappa)$ made any such calls prior to time $\tau$, then the two computations would have to diverge prior to the point where this happens for the first time, but we have just established, this cannot be. So neither computation makes a call to an ordinal in $\lambda,\kappa)$ up to time $\tau$, and the claim is proven.
Again, it is easy to see that this sequence of forcing is progressively closed, so the iteration works as in the proof of Theorem 3.3.
Question 1
This leaves us with at least two questions:
  1. 1.
    Can the large cardinal assumption be eliminated from the Theorem 3.4?
  2. 2.
    Is it consistent with ZFC that we have $\text{OrdCard}\nleq_{\text{OTM}}^{f}\text{DecCard}$ for any $f$ that satisfies $f(\alpha)<\alpha$ for all sufficiently large $\alpha$?
We conjecture that both answers are positive.

### v5

@@ 3.2 Reductions to DecCard
Remark 4
Considering the first call to some ordinal $\alpha\geq\kappa^{+}$ in the last lemma, we thus obtain that at least $\kappa^{+}$ many requests to $H$ have been made that concern ordinals strictly smaller than $\kappa^{+}$ prior to the first call to a larger ordinal.
We have already seen that the identity function is an upper bound on the reduction complexity of OrdCard to DecCard, as you can always start by testing the given ordinal $\alpha$ and then all ordinals below it successively. It is then natural to ask whether this is optimal: Is there (provably in ZFC) a reduction of OrdCard to DecCard which provably in ZFC improves on the naive approach explained above in the sense that, for some $\alpha\in\text{On}$, the number of calls required on input $\xi>\alpha$ will be strictly smaller than $\xi$? As the next theorem shows, this is not the case; it is thus consistent relative to ZFC that the upper bound id is in fact optimal. [^5] As a result, this is of course a strengthening of Theorem 3.3. However, the proof method here is much more specific to reductions to DecCard, which is why Theorem 3.3 – or rather its proof – is of independent interest.
@@
It is consistent relative to ZFC that any reduction of OrdCard to DecCard will, for unboundedly many $\alpha$, require at least $\alpha$ many calls to DecCard, i.e., that $\text{OrdCard}\nleq_{\text{OTM}}^{<\text{id}}\text{DecCard}$.
@@
As explained above, we identify parameter-programs with ordinals. For $\iota\in\text{On}$, we write $\kappa_{\iota}$ for the $\iota$-th fixed point of the $\aleph$ function in $L$. We will avoid complications for our construction that would arise from such fixed points that are themselves limits of such fixed points and thus only consider successors in this sequence; accordingly, we write $\kappa(\alpha,\beta)$ for $\kappa_{p(\alpha,\beta)+1}$. Thus, we have reserved, for each parameter-program $\alpha$, a proper class $C_{\alpha}$ of fixed points of the $\aleph$ function in $L$ (which are not limits of such fixed points), namely $C_{\alpha}:=\{\kappa(\alpha,\iota):\iota\in\text{On}\}$. We assume without loss of generality (e.g., by replacing $\kappa(\alpha,\beta)$ with $\kappa(\alpha,\beta+\alpha)$ if necessary) that $\kappa(\alpha,\beta)>\text{max}(\alpha,\beta)$.
@@
We now proceed by an iterated forcing, sabotaging, for each pair $(\gamma,\alpha)$ of a (coded) parameter-program $\gamma$ and an ordinal $\alpha$ the claim that, for all $\delta>\alpha$, $\gamma$ achieves a reduction of OrdCard to DecCard with strictly less than $\delta$ many calls to DecCard. Noting that DecCard is easily definable as a class function, let us denote by $H^{M}:\text{On}\rightarrow\{0,1\}$ an (in fact, the unique) DecCard function in the transitive ZFC-model $M$; when the model is clear from the context, the superscript $M$ will be dropped. By replacing the parameter-program $\gamma$ with the parameter-program $\gamma^{\prime}$ guaranteed to exist by Lemma 1, we can assume without loos of generality that each parameter-program $\gamma$ we consider calls $H$ at most once on each ordinal.
We now explain one step of the iteration, which will start with $L$. Let an ordered pair $(\gamma,\alpha)$ of a parameter-program $\gamma$ and an ordinal $\alpha$ be given. Consider the cardinal $\kappa:=\kappa(\gamma,\alpha)$. Since $\kappa$ is by definition not a limit of fixed points of the $\aleph$ function in $L$, there is a largest such fixed-point $\kappa^{\prime}$ below $\kappa$. Let $\delta\in((\kappa^{\prime})^{+},\kappa)$ be an $L$-cardinal, and let define the class function $\overline{H}$ as follows, for all $\xi\in\text{On}$:
$\overline{H}(\xi)=\begin{cases}0\text{, if }\delta<\xi\leq\kappa\\ H(\xi)\text{, otherwise.}\end{cases}$
We assume that our construction has already taken care of all earlier pairs in the natural well-ordering of pairs of ordinals and that we have thus obtained a transitive $M\models\text{ZFC}$ with the following properties: Denoting by $\nu$ the supremum of the fixed points of the $\aleph$-function below $\kappa$ in $L$ – which, by assumption, will be strictly smaller than $\nu$ – then, from $(\nu{+})^{M}$ upwards, the $M$-cardinals are the same as the $L$-cardinals. We will now pass to a generic extension $M[G]$ of $M$ with the following properties:
@@
    cardinals up to and including $(\nu^{+})^{M}$ are not changed
@@
    cardinals greater than or equal to $(\kappa^{+})^{M}$ are not changed (and thus remain the same as in $L$)
@@
    $\kappa$ is collapsed
  4. 4.
    for every cardinal $\mu\geq\kappa^{+}$, we have $2^{\mu}=\mu^{+}$ in $M[G]$
  5. 5.
    the computations $\gamma^{\overline{H}}(\kappa)$ and $\gamma^{H^{M[G]}}(\kappa)$ agree
  6. 6.
    either the computation $\gamma^{\overline{H}}(\kappa)$ makes at least $\kappa$ many calls to $\overline{H}$, or it does not halt with output $\text{card}^{M[G]}(\kappa)$.
Let $S$ be the set of ordinals to which the computation $\gamma^{\overline{H}}(\kappa)$ makes calls.
We distinguish the following cases:
@@
For some $\zeta<\kappa$, $(\zeta,\kappa)\cap S=(\zeta,\kappa)\cap\text{Card}^{L}$.
By assumption that $\kappa$ is a fixed-point of the $\aleph$ function, this implies in particular that $\gamma^{\overline{H}}(\kappa)$ makes at least $\kappa$ many calls to $\overline{H}$. In this case, $M[G]$ is obtained by collapsing $\kappa$ to $\delta$. This implies (1), (2), (3) and (4). Consequently, we have $\overline{H}=H^{M[G]}$ and thus (5). If this computation does not call $\overline{H}$ for any ordinal greater than or equal to $(\kappa^{+})^{M}$, then, because of (1), further extensions will not change it and we are done. If it does call $\overline{H}$ for some such ordinal, then, by Lemma 5, at least $(\kappa^{+})^{M}>\kappa$ many requests to $\overline{H}$ concerning ordinals strictly below $(\kappa^{+})^{M}$ have been made prior to it, and further extensions will not change the answers to these requests, so in this case as well, the computation $\gamma^{H}(\kappa)$ is guaranteed to make more than $\kappa$ many calls to $H$, which suffices. In both cases, we thus have the first option of (6).
@@
Otherwise.
Thus, there are cofinally in $\kappa$ many $L$-cardinals for which $\overline{H}$ is never called in the computation of $\gamma^{\overline{H}}(\kappa)$. Pick one that is greater than $\delta$ and different from the output of this computation (if there is one); call it $\lambda$.
Now, $M[G]$ is obtained by (i) collapsing $\kappa$ to $\lambda$ and (ii) collapsing all $L$-cardinals in $(\delta,\lambda)$ to $\delta$. Again, (1), (2), (3) and (4) are standard. To see (4), just note that the only argument for which $\overline{H}$ and $H^{M[G]}$ differ is $\lambda$, which, however, by our assumption, is not considered by this computation. Moreover, the construction guarantees that $\text{card}^{M[G]}(\kappa)=\lambda$ and that $\lambda$ is not the output of $\gamma^{H^{M[G]}}(\kappa)$, so we have the second alternative of (6).
In either case, we have (6): $\kappa$ is a counterexample to the claim that $\rho$ reduces OrdCard to DecCard on all instances above $\alpha$ in reduction complexity strictly less than id.
For the iteration, we again use Easton support. And again, since the iteration is progressively closed, it yields a model of ZFC in which $\text{OrdCard}\nleq_{\text{OTM}}{<\text{id}}\text{DecCard}$.
Remark 5
This does not mean that it is consistent that a reduction of OrdCard to DecCard will have to make calls to DecCard for all ordinals below the input; for example, it is clearly sufficient to only make requests for limit ordinals, or admissible ordinals, or admissible limits of admissible ordinals etc., and answer all requests to other ordinals with 0 without invoking the DecCard function.